OpenStudy (anonymous):

Provide an example that uses the substitution method to solve two linear equations.

OpenStudy (anonymous):

3x + y = 7 2x - 3y = 1 y = 7 - 3x 2x - 3(7 - 3x) = 1 2x - 21 - 9x = 1 11x = 22 x = 2 y = 7 - 3(2) = 7-6 = 1